Plss help me ans this question :a+3d/a+9d =a+d/a+5d=k, then find the value of k (a,d>0)

Grade:11

1 Answers

Vikas TU
7 years ago
a+3d/a+9d =a+d/a+5d
do cross multiyply,
(a+3d)(a+5d) = (a+d)(a+9d)
a^2 + 5ad + 3ad + 15d^2 = a^2 + 9ad + ad + 9d^2
-2ad + d^2 = 0
or
d^2 – 2ad = 0
d(d – 2a) = 0
d cannot be zero.
Therefore,
d = 2a
k = 3a/11a = 3/11.